Three 110.0-g ice cubes initially at 0°C are added to 0.900 kg of water initially at 21.0°C in an insulated container. (a) WhI am having trouble with part B and could use some help
Answer #1

I think you have done the problem correctly only you have not converted the answer into kg.

The ice is converted into water at 0°C decreasing the temperature of the 0.900 kg of water which is initially at 21°C. The amount of heat released by the water when its temperature changes from 21°C to 0°C is

Q=mC(21^oC-0^oC)

Q=(0.900\ kg)\times (4186\ J/(kg ^oC))\times(21^oC)

Q=79115\ J

The mass of ice that heat Q can melt is

M=\frac{Q}{L_{fusion}}

M=\frac{79115\ J}{334000\ J/kg}

M=0.2369\ kg

There were initially three ice cubes each of mass 110g. The initial mass of the ice is

M_o=3\times 110\ g=330\ g=0.330\ kg

The mass of unmelted ice is

\Delta M=M_o-M

\Delta M=0.330\ kg-0.2369\ kg

\boxed {\Delta M=0.0931\ kg}